OpenStudy (anonymous):

Find the point on the terminal side of θ = pi/4 that has an x coordinate of 2. Show your work for full credit. @mebs

OpenStudy (anonymous):

this is what I have so far when x=2, y=?z tanø= y/x

OpenStudy (anonymous):

alright pi/4 = 45 right... so what exact values are cos 0 and sin 0 for 45 degrees from the unit circle.

OpenStudy (anonymous):

sqrt2/2 and sqrt2/2?

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

yes the answer is 2... because tan 45 = 1 so if x/y = 1 than 2/2 =1